OpenStudy (anonymous):

You and a friend sign up for a health club. It costs $50 to join, plus $20 a month. However, the first 100 customers can sign up for $30, plus $20 a month. Your friend is customer 100, and you are 101. If the accumulated cost for each membership were graphed, what would it look like? What would the difference be between your accumulated cost and your friend's accumulated cost at any given month?

OpenStudy (anonymous):

Y = mx + c c is the cost to join . m is rate of change. your friend is y = 20x + 30 yours y = 20x + 50 the difference is 20 always.

OpenStudy (anonymous):

as catch me said... the graph would have a linear graph per y=20x + 30 until x=100, then the graph would continue as y=20x+50... since both have 20x, the slope is the same for both